12344444:

y = –x + 4 x + 2y = –8 How many solutions does this linear system have? one solution: (8, 0) one solution: (0, 8) no solution infinite number of solutions

Shadow:

\[x + 2(-x + 4) = -8\] \[x -2x + 8 = -8\] \[-x + 8 = -8\] \[x -8 = 8\] \[x = 16\] \[y = -16 +4 = -12\] x = 16 y = -12 (16, -12) That is one solution, but is not represented in your answer choices...